What percent of 144 is 126

Found 2 solutions by fcabanski, Alan3354:
Answer by fcabanski(1391)

The word of means multiply. Is means =


What percent of 144 in math terms: x% * 144
Is 126 in math terms: = 126.


x%*144=126
x%=126/144=.875 = 87.5%

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com


Answer by Alan3354(69443)

What percent of 144 is 126
---------
126/144 = 0.875
x100 --> 87.5%